a ray of light passes from air through a rectangular block of glass with parallel faces 6.5cm apart at an angle of incidence 35 degrees. calculate:
a) the angle of refraction
b) the lateral displacement
1
Expert's answer
2019-11-15T10:20:02-0500

Calculation of angle of refraction and the lateral displacement


We need to find the angle of refraction and lateral displacement


Answer:


Angle of incident = i = "35^o"


Angle of refraction = r


Refractive index of Air = "n_1 = 1"


Refractive index of glass with respect to air = "n_2 = 1.65"


thickness of the glass = t = 6.5cm


a).


Formula of Snell's Law



"\frac {sin i} {sin r} = \frac {n_2} {n_1}"



Now plug the values in the formula,


"\frac {sin \space 35^o} {sin r} = \frac {1.65} {1}"




"\frac {sin \space 35^o} {1.65} = \frac {sin r} {1}"


"{sin r} = \frac {sin \space 35^o} {1.65} = \frac {0.5735} {1.65} = 0.3475"

"r = \arcsin (0.3457) = 20.33^o"

(b).


Lateral displacement

"=\frac {t \times sin \space (i - r)} {cos r}"

Plug the values in the formula,



"Lateral \space displacement = \frac {t \times sin \space (i - r)} {cos r} =\frac { 6.5 \times sin (35^o - 20.33^o)} {cos \space 20.33^o}"

=

"= \frac {6.5 \times 0.8613} {0.09} =62.2 cm"